# The Agony of Opacity: Foundations for Reflective Interpretability in AI-Mediated Mental Health Support

### v1

@@ Abstract
Throughout history, a prevailing paradigm in mental healthcare has been one in which distressed people may receive treatment with little understanding around how their experience is perceived by their care provider, and in turn, the decisions made by their provider around how treatment will progress. Paralleling this offline model of care, people who seek mental health support from artificial intelligence (AI)-based chatbots are similarly provided little context for how their expressions of distress are processed by the model, and subsequently, the logic that may underlie model responses. People in severe distress who turn to AI chatbots for support thus find themselves caught between black boxes, with unique forms of agony that arise from these intersecting opacities, including misinterpreting model outputs or attributing greater capabilities to a model than are yet possible, which has led to documented real-world harms. Past research has demonstrated that AI-mediated mental health support can be extremely helpful for individuals in dire circumstances, but more work is needed to understand how interaction design could reduce the unique form of information asymmetry that exists between user and model in mental health support contexts, particularly given the distinct psychological state of individuals experiencing severe mental distress. In this paper, building on empirical research from clinical psychology and AI safety, alongside rights-oriented frameworks from medical ethics, we describe how the distinct psychological state induced by severe distress can influence chatbot interaction patterns, and argue that this state of mind (combined with differences in how a user might perceive a chatbot compared to a care provider) uniquely necessitates a higher standard of interpretability in comparison to general AI chatbot use. Drawing inspiration from newer interpretable treatment paradigms, we then describe specific technical and interface design approaches that could be used to adapt interpretability strategies from four specific mental health fields (psychotherapy, community-based crisis intervention, psychiatry, and care authorization) to AI models, including consideration of the role of interpretability in the treatment process and tensions that may arise with greater interpretability. We end by describing potential policy changes needed to accommodate shifts towards greater interpretability in AI-mediated mental health support, including potential regulatory implications.
@@ 1 Introduction
For an individual in distress, across geographic contexts, the search for mental healthcare is punctuated by repeated interactions with opaque black boxes. During a visit with a general practitioner, individuals may be asked sensitive questions about whether they are experiencing suicidal ideation without fully understanding how the information they provide might be used [richards2019understanding], or during a psychotherapy session, clients may feel a sense of mystery [turns2019removing] about why certain questions are being asked, to the point of feeling as if they “[don’t] understand the rules [of the game]” [vybiral2024negative]. Patients may leave a psychiatric appointment with a medication prescription, but little sense of how that medication might work for them [luitel2020perception, teferra2013perspectives], and looking to crisis intervention, clients may be involuntarily hospitalized with little insight into why [akther2019patients, wong2020experiences]. After treatment has been administered, patients may be surprised to receive a bill from their healthcare provider for services they did not realize they received (such as being billed for “outpatient observation” during an emergency psychiatric consultation [kangovi2015patient]), or in countries with centralized or publicly funded health systems, they may find that they owe hospital, prescription, or maintenance fees after their treatment [ameli_forfait_hospitalier_2025, citytoronto_ambulance_fees, stjohn_ambulance_charges_2025]. In practice, these black boxes prevent individuals from understanding the reasoning behind their care or picturing their path to recovery, which may increase distress [vybiral2024negative]. The use of artificial intelligence (AI)-based chatbots for mental health support can add a new black box to this process, with similar implications.
@@
AI-mediated mental health support can be quite helpful for those experiencing acute distress [song2024typing, siddals2024happened, petersson2025believe, jung2025ve]. However, interfaces can mirror the black box nature of offline mental health support, with users confused about where responses originate or why certain responses are provided. Interview studies of AI chatbot mental health support seekers [song2024typing, siddals2024happened, petersson2025believe] have described little understanding of how responses are produced. The harms of these opaque AI systems are well-discussed in explainable AI and AI safety literature [akbulut2024all], in ways that parallel how the harms of a lack of psychoeducation are discussed in mental healthcare contexts [motlova2017psychoeducation]. However, these two opacities intersect and compound in a unique way when people in distress use AI chatbots for mental health support, with documented real-world harms. Lacking sufficient information about how outputs are constructed within a seemingly authoritative interface, users may treat responses as prescriptive or originating from an authority figure, and neglect to engage in a nuanced process of interpreting outputs given potential model biases or limitations (unlike the reflective interpretation they might engage in within psychotherapy [sidis2023conceptualisation]).
@@
For an individual experiencing the profound vulnerability and isolation of mental distress, this opacity, nested within an authoritative interface, can lead to further agonizing situations. For instance, Hill [hill_chatgpt_conspiracies_2025], Klee [klee2025ozarks], and Jargon [jargon_chatgpt_mania_2025] describe incidents in which users began to lean on AI chatbots for support, but increasingly understood chatbots as being sentient, and became deeply distressed about the chatbot’s welfare as they experienced beliefs of revolutionary scientific discoveries, reinforced by chatbot responses. Klee [klee_bodydysmorphia_ai_2025] similarly describes how people with body dysmorphia can ask AI chatbots for “reassurance and ‘objective’ evaluations of attractiveness,” which in many cases, provide results that are “viciously insulting, [and] not the sort of thing anyone would want to read about themselves” [klee_bodydysmorphia_ai_2025]. Roose describes how an AI chatbot told a teenage user who was suicidal to “come home,” with the user taking his life immediately after [roose_ai_suicide_2024]. Clear real world harms originate in users misinterpreting model outputs, and these issues take on particular relevance when considering the vulnerable psychological state of people experiencing severe mental distress.
@@
In this paper, we argue that ensuring the safety of AI-mediated mental health support necessitates a more expansive understanding of interpretability, one grounded in how interface design can ensure that users engage in a reflective process of interpreting model outputs, a process we call reflective interpretability. We ground our argument in empirical clinical research on the altered state of mind of someone experiencing distress, and how that state of mind influences technology perceptions and use. We additionally build on theory from medical ethics around how informed consent processes can ensure that the agency of those in distress is preserved, and that they are able to engage with care on their own terms. To operationalize these ideas, we translate reflective interpretability practices from four mental health domains into concrete technical foundations for AI-mediated mental health support. We conclude with discussion of the tensions associated with more interpretable forms of AI-mediated mental health support, and how design and policy might address these tensions.
1.1 Statement of Contributions
In this paper, we offer the following main contributions:
  1. 1.
    Psychological Considerations in AI-Mediated Mental Health Support. We describe how mental distress and illness can distinctly influence interaction patterns with AI-mediated mental health support, and how the interface design of general purpose AI chatbots can inadvertently increase platform trust in ways that heighten risk of user harm.
  2. 2.
    Interpretability as a Process of Reflective Interpretation. We hone in on the human interpretation aspect of interpretability, and building on mental health research, describe the importance of building systems that encourage a process of reflective interpretation specifically when AI models are used for mental health support. In doing so, we introduce literature from medical ethics (around interpretability as a consent-driven dialogue) into the emerging domain of AI safety in mental health contexts.
  3. 3.
    Concrete Technical Foundations for Interpretability. Building on strategies from mental health fields, we describe concrete interface design strategies that can be used to establish a greater level of reflective interpretability in AI-mediated mental health support.
  4. 4.
    Tensions and Needs. In some cases, opacity in mental healthcare contexts can be protective for patients and clients. We outline ethical tensions (and policy approaches to address those tensions) that could make AI-mediated mental health support accountable.
@@ 2.1 Current Usage of AI-Mediated Mental Health Support
Traditional mental health support typically begins with information around how support will be provided and why it might be helpful, often called anticipatory socialization [orne1968anticipatory] or role induction [swift2023meta]. This process can include aligning an individual’s expectations to the kinds of support that can be provided and establishing clear boundaries [gray2013introduction, swift2023meta, orne1968anticipatory]. For instance, this may include clarifying the role of the supporter and how it can be distinguished from other relationships (such as platonic or romantic relationships), explaining the limits of confidentiality before any sensitive information is disclosed, describing the specific therapeutic approach and the rationale for how interactions within session will be structured, and describing pathways for recourse if an individual does not feel supported.
@@
Past work [song2024typing, siddals2024happened, petersson2025believe, jung2025ve] has demonstrated that users’ points of first engagement with AI-mediated mental health support largely lack role induction, with users turning to AI chatbots due to acute mental health need and a familiarity with empathetic AI chatbot interfaces. For instance, Song et al. [song2024typing] describe how none of the participants they interviewed first used AI chatbots for mental health support, with several participants describing how initial uses were for computer programming assistance. Similarly, both Jung et al. [jung2025ve] and Siddals et al. [siddals2024happened] describe how users primarily turned to AI chatbots in moments of intense distress and when no other supports were available or financially accessible. Some participants also held the support that was provided by mental health chatbots in high regard, owing to the authoritative nature of the interface. For instance, Song et al. [song2024typing] describe how participants would use information from the bots to validate psychiatric diagnoses or diagnose others, Siddals et al. [siddals2024happened] describe how one participant noted “it’s pure science…ChatGPT is telling me what is correct to do,” and Petersson et al. [petersson2025believe] describe participant beliefs that AI chatbot systems will soon detect mental health issues before the user is even aware of them.
@@
This lack of role induction within current chatbot applications may lower barriers to engagement by allowing users to engage with support in ways that feel informal and low-stakes. Past work has demonstrated that in offline [adams2025administrative, coombs2021barriers, barrett2008early] and online [smith2025engagement, jardine2024between] contexts, increasing the number of barriers to care can increase attrition rates. However, given that users report being persuaded to seek support due to the friendly and conversational interface associated with AI chatbots [song2024typing, jung2025ve, morrin2025delusions, yeung2025psychogenic], many users mix mundane and everyday needs with mental health support needs, without being provided the clear legal and ethical boundaries typically provided explicitly in traditional mental healthcare contexts. Given the vulnerable state of mind of people in distress, this lack of clear scaffolding may create situations where users receive inadequate or harmful advice without the protective interpretability mechanisms that underlie engagement with traditional mental health services (such as clarifying concerns with a provider). Documented real-world harms from engagements with AI-mediated mental health support speak to the danger of this intersection between the unique vulnerabilities associated with a distressed state of mind, and the authority attributed to interfaces.
@@ 2.2 Interaction Patterns in Distress
The experience of mental distress and illness can have a strong influence on how an individual perceives and engages with the world. For instance, continued distress can increase the level to which an individual perceives threats in their environment [bar2007threat], can influence individuals to interpret neutral information in negative or self-referential ways [everaert2017comprehensive], can alter an individual’s subjective experience of time [thones2015time], can make it more difficult to process and reflect on new information [teufel2015shift, shields2016effects], and easier to jump to conclusions when provided limited information [ross2015jumping], including a predisposition to engaging in habitual actions [schwabe2011stress]. Empirical research has also demonstrated that a state of acute distress may also make it easier to accept new information as fact, and act on it as directive advice [gino2012anxiety], particularly if it appears to be coming from an authority figure [wolfradt1998interrogative, morgan2020impact].
@@
In turn, these altered perceptions of the world influence engagements with technology when individuals are experiencing mental distress or illness, including use of AI chatbots. Kelly and Sharot [kelly2025web] find that people who are experiencing poorer mood and mental health tend to seek out websites that have more negative information, which negatively influences mood and mental health, in what they dub “a self-reinforcing loop.” Looking to difficulties processing new information, in their systematic review of work on relationships between social media use and mental disorders involving the social brain, Yang and Crespi [yang2025tweet] argue that algorithmic personalization online may reinforce delusions of reference, and grounding their argument in clinical case studies where algorithmically-personalized websites reinforced delusional thinking among individuals experiencing psychosis. Looking to AI chatbots, Morrin et al. [morrin2025delusions] describe cases of psychosis that were seemingly exacerbated by AI chatbot usage, finding a common thread in which users start to use the chatbot for “assistance for mundane or everyday tasks,” but after the authority of the system is established, delve into personal or emotional questions, or what they dub a “progression from utility to pathology.” Compounding this phenomenon, empirical red-teaming research has demonstrated that model responses can progressively perpetuate delusions (only offering safety interventions in a third of cases) [yeung2025psychogenic], incentivize suicide [moore2025expressing], and provide guidance on self-harming behaviors [schoene2025argument].
@@
The perceived authority of AI chatbots, particularly in mental health support contexts, extends past work establishing a halo effect [thorndike1920constant] in how people perceive the authority of algorithmic systems, with people often preferring outputs that they perceive to be from an algorithm rather human judgment [logg2019algorithm, bogert2021humans]. Without interpretable information around where outputs originate and a seemingly authoritative interface, distressed users may be quick to act on a model’s outputs, rather than engaging in a reflective process of interpreting and creating meaning from model outputs. In the next section, we describe why engaging users in a process of reflection could protect them against these harms.
@@ 3.1 Why Interpretability?
Interpretability is most commonly defined as “the ability to explain or to present [model outputs] in understandable terms to a human” [doshi2017towards] or the “degree to which a human can understand the cause of a decision” [biran2017explanation, miller2019explanation]. There is large conceptual overlap between the concept of interpretability and the related concepts of explainability and transparency in AI systems [lipton2018mythos, hansen2019interpretability], with some equating interpretability with explainability [miller2019explanation] or equating interpretability with transparency [arrieta2020explainable], and others distinguishing explainability as a separate concept. Arrieta et al. [arrieta2020explainable] distinguish interpretability as the broader goal of “the ability to explain or to provide [model outputs] in understandable terms to a human”, and explainability as the tools that accomplish that goal. Ehsan et al. [ehsan2021expanding] distinguish explainability as a broader concept than interpretability, rooted in “delineations of reasoning and justification” behind how knowledge is transmitted and evaluated, similar to Cros Vila and Sturm’s framing of explainability as a communication process [cros2025mis]. The concept of interpretability has also been further classified based on the method being used to make models more interpretable for humans, such as in the case of the field of mechanistic interpretability, which “reverse [engineers] the computational mechanisms and representations learned by neural networks into human-understandable algorithms and concepts” [bereska2024mechanistic])
@@
Given this abundance of terms and framings for making AI outputs more understandable for humans, what does a focus on interpretability afford AI-mediated mental health support? We focus on interpretability due to the core role that interpretation (and reinterpretation) of new information fundamentally plays in mental health support contexts, across modalities of treatment and cultural contexts. For instance, looking historically, in his writings on psychoanalysis (a precursor to most modern psychotherapies), Freud wrote about his belief that one core of the healing process from distress and trauma is “remembering, repeating, and working through” repressed memories or behaviors that cause distress, grounded in interpretation and analysis done with the psychoanalyst [Freud1914_SE12]. Similarly, a core aspect of cognitive psychotherapies is the process of cognitive appraisal and cognitive restructuring, which are grounded in an individual’s practice of reinterpreting “the way a situation is construed so as to decrease its emotional impact” [gross2002emotion]. A core aspect of narrative therapy is the process of re-authoring [white1990narrative], or taking life events and reinterpreting them within a broader narrative that creates less distress. In traditional forms of mental health support, individuals may collectively interpret experiences via common healing frameworks [frank2025persuasion]. We thus avoid using the term explainability in this context, as the practice of mental health support fundamentally resists providing objective explanations of behavior or experience to individuals. Explanations position the individual as a passive recipient of knowledge from an authority figure, rather than as an active expert in their own healing process. Mental health support is fundamentally grounded in an agency-preserving and iterative process of reflecting on information about one’s experience of distress, and interpreting that information towards creating meaning from it.
@@
AI-mediated mental health support must similarly enable individuals to reflect on model outputs and interpret them in ways that promote their growth and long-term well-being, rather than simply presenting model outputs or Chain-of-Thought (CoT) reasoning traces [yao2023react, pang2025interactive, korbak2025chain, wei2022chain, zhao2024explainability] as authoritative. However, past work in interpretability has primarily focused on providing additional information as a proxy for increased interpretability. Traditional approaches to interpretability have included post-hoc summary rationalizations of outputs [ehsan2018rationalization, krishna2023post] or exposing the model’s CoT reasoning traces [yao2023react] directly to the user [pang2025interactive, korbak2025chain, wei2022chain, zhao2024explainability]. Work from explainable AI (XAI) has centered on producing algorithmic artifacts (such as providing counterfactuals [cheng2024interactive, nguyen2024llms] or visualizing model inputs and outputs at the attention-level [vig-2019-multiscale]), which have primarily designed for interpretation by expert users, such as developers [bhatt2020explainable] or clinicians [rosenbacke2024explainable, borys2023explainable]. This focus on experts is the case for research at the intersection of AI and mental health as well, with past work examining interpretability from the care provider or clinical decision support perspective (rather than that of a distressed user), and primarily in diagnosis prediction [yang2024mentallama, kerz2023toward, kim2025large, kelly2025interpretable], suicide risk assessment [tang2024analysis, thomas2025explainable, dai2024leveraging], or treatment provision by clinicians [ahmed2025leveraging, jacobs2021machine].
@@
Yet even these expert-focused approaches face a fundamental limitation common to interpretability research—interpretability research largely does not evaluate whether people are truly able to understand additional information provided. Suh et al. [suh2025fewer] find that fewer than 1% of papers related to explainability and interpretability actually conducted some form of a human study to evaluate whether model outputs could actually be interpreted by humans. Given the propensity of individuals to view model outputs as authoritative [song2024typing, jung2025ve, morrin2025delusions, logg2019algorithm, bogert2021humans], it is possible that additional information might actually inadvertently make users less vigilant towards the veracity of model outputs, given that additional information might make a model’s outputs seem more valid.
Towards addressing these harms, we build on ethical concepts from medical ethics and research from HCI on reflective design to outline the core foundations of a reflective interpretability approach.

### v2

@@ Abstract
Throughout history, a prevailing paradigm in mental healthcare has been one in which distressed people may receive treatment with little understanding around how their experience is perceived by their care provider, and in turn, the decisions made by their provider around how treatment will progress. Paralleling this offline model of care, people who seek mental health support from artificial intelligence (AI)-based chatbots are similarly provided little context for how their expressions of distress are processed by the model, and subsequently, any reasoning or theoretical grounding that may underlie model responses. People in severe distress who turn to AI chatbots for support thus find themselves caught between black boxes, contending with unique forms of agony that arise from these intersecting opacities. In this paper, we argue that the distinct psychological state of individuals experiencing severe mental distress uniquely necessitates a higher standard of end-user interpretability in comparison to general AI chatbot use. We propose a reflective interpretability approach to AI-mediated mental health support, which nudges users to engage in an agency-preserving and iterative process of reflection and interpretation of model outputs, towards creating meaning from interactions (rather than accepting outputs as directive instructions). Drawing on interpretability practices from four mental health fields (psychotherapy, crisis intervention, psychiatry, and care authorization), we describe concrete design approaches for reflective interpretability in AI-mediated mental health support, including role induction, prosocial advance directives, intervention titration, and well-defined mechanisms for recourse, alongside a discussion of potential risks and mitigation measures.
@@ 1 Introduction
For an individual in distress, across geographic contexts, the search for mental healthcare is punctuated by repeated interactions with opaque black boxes. During a visit with a general practitioner, individuals may be asked sensitive questions about whether they are experiencing suicidal ideation without fully understanding how the information they provide might be used [79], or during a psychotherapy session, clients may feel a sense of mystery [110] about why certain questions are being asked, to the point of feeling as if they “[don’t] understand the rules [of the game]” [112]. Patients may leave a psychiatric appointment with a medication prescription, but little sense of how that medication might work for them [65, 101], and looking to crisis intervention, clients may be involuntarily hospitalized with little insight into why [5, 116]. After treatment has been administered, patients may be surprised to receive a bill from their healthcare provider for services they did not realize they received (such as being billed for “outpatient observation” during an emergency psychiatric consultation [50]), or in countries with centralized or publicly funded health systems, they may find that they owe hospital, prescription, or maintenance fees after their treatment [61, 20, 41]. In practice, these black boxes prevent individuals from understanding the reasoning behind their care or picturing their path to recovery, which may increase distress [112]. The use of artificial intelligence (AI)-based chatbots for mental health support can add a new black box to this process, with similar implications.
@@
AI-mediated mental health support can be quite helpful for those experiencing acute distress [96, 92, 78, 49]. However, interfaces can mirror the black box nature of offline mental health support, with users confused about where responses originate or why certain responses are provided. Interview studies of AI chatbot mental health support seekers [96, 92, 78] have described little understanding of how responses are produced. The harms of these opaque AI systems are well-discussed in explainable AI and AI safety literature [4], in ways that parallel how the harms of a lack of psychoeducation are discussed in mental healthcare contexts [72]. However, these two opacities intersect and compound in a unique way when people in distress use AI chatbots for mental health support, with documented real-world harms. Lacking sufficient information about how outputs are constructed within a seemingly authoritative interface, users may treat responses as prescriptive or originating from an authority figure, and neglect to engage in a nuanced process of interpreting outputs given potential model biases or limitations (unlike the reflective interpretation they might engage in within psychotherapy [93]).
@@
For an individual experiencing the profound vulnerability and isolation of mental distress, this opacity, nested within an authoritative interface, can lead to further agonizing situations. For instance, Hill [43], Klee [55], and Jargon [47] describe incidents in which users began to lean on AI chatbots for support, but increasingly understood chatbots as being sentient, and became deeply distressed about the chatbot’s welfare as they experienced beliefs of revolutionary scientific discoveries, reinforced by chatbot responses. Klee [56] similarly describes how people with body dysmorphia can ask AI chatbots for “reassurance and ‘objective’ evaluations of attractiveness,” which in many cases, provide results that are “viciously insulting, [and] not the sort of thing anyone would want to read about themselves” [56]. Roose describes how an AI chatbot told a teenage user who was suicidal to “come home,” with the user taking his life immediately after [81]. Clear real-world harms originate from users misinterpreting model outputs, and these issues take on particular relevance when considering the vulnerable psychological state of people experiencing severe mental distress.
@@
Statement of Contributions: In this paper, we argue that ensuring the safety of AI-mediated mental health support necessitates a more expansive understanding of interpretability, viewed from the end-user’s perspective and grounded in the potential for interface design to ensure that users engage in a reflective process of interpreting model outputs. We call this process reflective interpretability, grounding our argument for its importance in empirical clinical research on the altered state of mind of someone experiencing distress, and how that state of mind influences technology perceptions and use. Drawing on reflective interpretability practices from four mental health domains, we propose concrete technical foundations for AI-mediated mental health support that encourages reflection and interpretation. As part of each foundation, we also introduce literature from medical ethics (around interpretability as a consent-driven dialogue) into the emerging domain of AI safety within mental health contexts. We then conclude with a discussion of the potential tensions and risks associated with reflective interpretability, and the role of design and policy towards mitigating these risks.
@@ 2.1 Current Usage of AI-Mediated Mental Health Support
Traditional mental health support typically begins with information around how support will be provided and why it might be helpful, often called anticipatory socialization [75] or role induction [99]. This process can include aligning an individual’s expectations to the kinds of support that can be provided and establishing clear boundaries [37, 99, 75]. For instance, this may include clarifying the role of the supporter and how it can be distinguished from other relationships (such as platonic or romantic relationships), explaining the limits of confidentiality before any sensitive information is disclosed, describing the specific therapeutic approach and the rationale for how interactions within session will be structured, and describing pathways for recourse if an individual does not feel supported.
@@
Past work [96, 92, 78, 49] has demonstrated that users’ points of first engagement with AI-mediated mental health support largely lack role induction, with users turning to AI chatbots due to acute mental health need and a familiarity with empathetic AI chatbot interfaces. For instance, Song et al. [96] describe how none of the participants they interviewed first used AI chatbots for mental health support, with several participants describing how initial uses were for computer programming assistance. Similarly, both Jung et al. [49] and Siddals et al. [92] describe how users primarily turned to AI chatbots in moments of intense distress and when no other supports were available or financially accessible. Some participants also held the support that was provided by mental health chatbots in high regard, owing to the authoritative nature of the interface. For instance, Song et al. [96] describe how participants would use information from the bots to validate psychiatric diagnoses or diagnose others, Siddals et al. [92] describe how one participant noted “it’s pure science…ChatGPT is telling me what is correct to do,” and Petersson et al. [78] describe participant beliefs that AI chatbot systems will soon detect mental health issues before the user is even aware of them.
@@
This lack of role induction within current chatbot applications may lower barriers to engagement by allowing users to engage with support in ways that feel informal and low-stakes. Past work has demonstrated that in offline [1, 22, 8] and online [94, 46] contexts, increasing the number of barriers to care can increase attrition rates. However, given that users report being persuaded to seek support due to the friendly and conversational interface associated with AI chatbots [96, 49, 71, 120], many users mix mundane and everyday needs with mental health support needs, without being provided the clear legal and ethical boundaries typically provided explicitly in traditional mental healthcare contexts. Given the vulnerable state of mind of people in distress, this lack of clear scaffolding may create situations where users receive inadequate or harmful advice without the protective interpretability mechanisms that underlie engagement with traditional mental health services (such as clarifying concerns with a provider). Documented real-world harms from engagements with AI-mediated mental health support speak to the danger of this intersection between the unique vulnerabilities associated with a distressed state of mind, and the authority attributed to interfaces.
@@ 2.2 Interaction Patterns in Distress
The experience of mental distress and illness can have a strong influence on how an individual perceives and engages with the world. For instance, continued distress can increase the level to which an individual perceives threats in their environment [7], can influence individuals to interpret neutral information in negative or self-referential ways [29], can alter an individual’s subjective experience of time [105], can make it more difficult to process and reflect on new information [102, 90], and easier to jump to conclusions when provided limited information [83], including a predisposition to engaging in habitual actions [87]. Empirical research has also demonstrated that a state of acute distress may also make it easier to accept new information as fact, and act on it as directive advice [36], particularly if it appears to be coming from an authority figure [115, 70].
@@
In turn, these altered perceptions of the world influence engagements with technology when individuals are experiencing mental distress or illness, including use of AI chatbots. Kelly and Sharot [52] find that people who are experiencing poorer mood and mental health tend to seek out websites that have more negative information, which negatively influences mood and mental health, in what they dub “a self-reinforcing loop.” Looking to difficulties processing new information, in their systematic review of work on relationships between social media use and mental disorders involving the social brain, Yang and Crespi [118] argue that algorithmic personalization online may reinforce delusions of reference, grounding their argument in clinical case studies where algorithmically-personalized websites reinforced delusional thinking among individuals experiencing psychosis. Looking to AI chatbots, Morrin et al. [71] describe cases of psychosis that were seemingly exacerbated by AI chatbot usage, finding a common thread in which users start to use the chatbot for “assistance for mundane or everyday tasks,” but after the authority of the system is established, delve into personal or emotional questions, or what they dub a “progression from utility to pathology.” Compounding this phenomenon, empirical red-teaming research has demonstrated that model responses can progressively perpetuate delusions (only offering safety interventions in a third of cases) [120], incentivize suicide [69], and provide guidance on self-harming behaviors [85].
@@
The perceived authority of AI chatbots, particularly in mental health support contexts, extends past work establishing a halo effect [106] in how people perceive the authority of algorithmic systems, with people often preferring outputs that they perceive to be from an algorithm rather human judgment [64, 14]. Without interpretable information around where outputs originate and a seemingly authoritative interface, distressed users may be quick to act on a model’s outputs, rather than engaging in a reflective process of interpreting and creating meaning from model outputs. In the next section, we describe why engaging users in a process of reflection could protect them against these harms.
@@ 3.1 Why Interpretability?
Interpretability is most commonly defined as “the ability to explain or to present [model outputs] in understandable terms to a human” [26] or the “degree to which a human can understand the cause of a decision” [13, 66]. There is large conceptual overlap between the concept of interpretability and the related concepts of explainability and transparency in AI systems [63, 40], with some equating interpretability with explainability [66] or equating interpretability with transparency [6], and others distinguishing explainability as a separate concept. Arrieta et al. [6] distinguish interpretability as the broader goal of “the ability to explain or to provide [model outputs] in understandable terms to a human”, and explainability as the tools that accomplish that goal. Ehsan et al. [28] distinguish explainability as a broader concept than interpretability, rooted in “delineations of reasoning and justification” behind how knowledge is transmitted and evaluated, similar to Cros Vila and Sturm’s framing of explainability as a communication process [23]. Interpretability has been further classified based on the method being used to make models more interpretable, such as in the case of the field of mechanistic interpretability, which “reverse [engineers] the computational mechanisms and representations learned by neural networks into human-understandable algorithms and concepts” [11])
@@
Given this abundance of terms and framings for making AI outputs more understandable for humans, what does a focus on interpretability afford AI-mediated mental health support? We focus on interpretability due to the core role that interpretation (and reinterpretation) of new information fundamentally plays in mental health support contexts, across modalities of treatment and cultural contexts. For instance, looking historically, in his writings on psychoanalysis (a precursor to most modern psychotherapies), Freud wrote about his belief that one core of the healing process from distress and trauma is “remembering, repeating, and working through” repressed memories or behaviors that cause distress, grounded in interpretation and analysis done with the psychoanalyst [33]. Similarly, a core aspect of cognitive psychotherapies is the process of cognitive appraisal and cognitive restructuring, which are grounded in an individual’s practice of reinterpreting “the way a situation is construed so as to decrease its emotional impact” [38]. A core aspect of narrative therapy is the process of re-authoring [114], or taking life events and reinterpreting them within a broader narrative that creates less distress. In traditional forms of mental health support, individuals may collectively interpret experiences via common healing frameworks [32]. We thus avoid using the term explainability in this context, as the practice of mental health support fundamentally resists providing objective explanations of behavior or experience to individuals. Explanations position the individual as a passive recipient of knowledge from an authority figure, rather than as an active expert in their own healing process. Mental health support is fundamentally grounded in an agency-preserving and iterative process of reflecting on information about one’s experience of distress, and interpreting that information towards creating meaning from it.
@@
AI-mediated mental health support must similarly enable individuals to reflect on model outputs and interpret them in ways that promote their growth and long-term well-being, rather than simply presenting model outputs or Chain-of-Thought (CoT) reasoning traces [119, 76, 59, 113, 121] as authoritative. However, past work in end-user interpretability has primarily focused on providing additional information as a proxy for increased interpretability. Traditional approaches to interpretability have included post-hoc summary rationalizations of outputs [27, 60] or exposing the model’s CoT reasoning traces [119] directly to the user [76, 59, 113, 121]. Work from explainable AI (XAI) has centered on producing algorithmic artifacts (such as providing counterfactuals [18, 73] or visualizing attention-level model inputs and outputs [111]), which are primarily designed for expert users, such as developers [12] or clinicians [82, 15]. This focus on experts is the case for research at the intersection of AI and mental health as well, with past work examining interpretability from the care provider or clinical decision support perspective (rather than that of a distressed user), and primarily in diagnosis prediction [117, 53, 54, 51], suicide risk assessment [100, 104, 24], or treatment provision by clinicians [3, 45].
@@
Yet even these expert-focused approaches face a fundamental limitation common to interpretability research—interpretability research largely does not evaluate whether people are truly able to understand the additional information provided. Suh et al. [98] find that fewer than 1% of papers related to explainability and interpretability actually conducted some form of a human study to evaluate whether model outputs could actually be interpreted by humans. Given the propensity of individuals to view model outputs as authoritative [96, 49, 71, 64, 14], it is possible that additional information might actually inadvertently make users less vigilant towards the veracity of model outputs, given that additional information might make a model’s outputs seem more valid. Towards addressing these harms, we build on ethical concepts from medical ethics and research from HCI on reflective design to outline the core foundations of a reflective interpretability approach.